OpenStudy (anonymous):

The effective rate of a $25,000 non-interest-bearing simple discount 10%, 90-day note is A. 10.62%. B. 10.8%. C. 10.26%. D. 10%.

OpenStudy (tkhunny):

25000 due in 90 days. I'll use 365 days per year. 10% simple discount: 25000*0.10(90/365) = 616.44 Cash in hand at the beginning of the 90 days: 25000 - 616.44 = 24,383.56 Solve for r: 616.44 = 24383.56*r*(90/365) ==> r = 0.10252837 I might also try 360 days. There may be pertinent classroom discussion on this.